OpenStudy (karatechopper):

use the points (1998,13.60) and (2003,19.00) to write the slope intercept form into an equation. plz guide me through this. i got a negative and i dont believe its right show me ur steps use the points (1998,13.60) and (2003,19.00) to write the slope intercept form into an equation. plz guide me through this. i got a negative and i dont believe its right show me ur steps @Mathematics

hero (hero):

Find the slope

OpenStudy (karatechopper):

i kno what to do but can u show me the steps?

OpenStudy (anonymous):

slope = 5.4/5 y=5.4/5(x-2003)+19 ... and you can simplify this

hero (hero):

19.00 - 13.60/2003 - 1998

OpenStudy (karatechopper):

look, the slope is 1.08 but i need help on the rest

OpenStudy (anonymous):

y = slope(x- a) + b ... the point is (a,b)

OpenStudy (anonymous):

so y = 1.08(x-2003) + 19

OpenStudy (anonymous):

you can expand and simpify this right??

OpenStudy (karatechopper):

i could but i wanna see if i got it right

OpenStudy (anonymous):

y=1.08x -1.08(2003) +19 y=1.08x -2144.24

OpenStudy (karatechopper):

so i got it right!! yes!!!